OpenStudy (anonymous):

statistics question. will medal and all that good stuff. Emily is playing a board game that has a spinner divided into equal sections numbered 1 to 18. The probability of the spinner landing on an even number or a multiple of 3 is BLANK. The probability of the spinner landing on an odd number or a number between 4 and 15 (both numbers excluded) is BLANK.

OpenStudy (anonymous):

first blank options; 1/6, 1/3, 1/2, 2/3, or 5/6

OpenStudy (anonymous):

second blank options; are the same numbers to choose from

OpenStudy (reemii):

The general method is the most boring but always works. Write down all numbers from 1 to 18. For each of them, check if it verifies the property described ("even number OR multiple of 3"). Count the numbers that verify this property: \(k\). Answer is \(\frac{k}{18}\).

OpenStudy (anonymous):

bless ur soul lol thank you so much

OpenStudy (reemii):

Yes :D For first question, No: 1, 5, 7, 11, 13, 17 Yes, 2, 3, 4, 6, 8, 9, 10, 12, 14, 15, 16, 18. Answer: \(\frac{12}{18}=\frac23\).

OpenStudy (reemii):

Yw* lol

OpenStudy (reemii):

What do you find for the second question?

OpenStudy (anonymous):

there is no second question just a second blank which is the second sentence that is stated

OpenStudy (reemii):

Yeah, that's what I meant. Did you solve it?

OpenStudy (anonymous):

no

OpenStudy (reemii):

Please try, we'll share answers :-)

OpenStudy (anonymous):

1/3?

OpenStudy (reemii):

It should be higher. Numbers between 4 and 15 are: 5,6,7,8,9,10,11,12,13,14. (10 numbers) Add the ones from (1,2,3,4,15,16,17,18) which are odd.

OpenStudy (reemii):

I count 4 numbers. Total: 14 numbers. Answer 14/18. But 7/9 isn't in the options. Where did it go wrong?

OpenStudy (reemii):

I don't know. Good luck.

OpenStudy (anonymous):

no idea. I really appreciate the help though thank you

OpenStudy (dinah_rose):

1,3,5,6,7,8,9,10,11,12,13,14,15,16,17 15/18 = 5/6

OpenStudy (reemii):

I don't think 16 is odd or between 4 and 15.

OpenStudy (dinah_rose):

ohh i thought it said 18 sorry

OpenStudy (anonymous):

I appreciate the help. I just used my best guess so I guess we'll see. Thanks :-)

OpenStudy (dinah_rose):

let me know the answer so i can see what went wrong n how we got 14/18 = 7/9

OpenStudy (anonymous):

Unfortunately plato does not EVER tell me what I get wrong so I fail to learn. It sucks. :-((